# CSDO: Enhancing Efficiency and Success in Large-Scale Multi-Vehicle Trajectory Planning

### v1

@@ IV-A Simulation Settings
The benchmark consists of various map sets with different sizes (50m $\times$ 50m, 100m $\times$ 100m), varying numbers of agents (5 to 100), and includes both obstructed and obstacle-free maps. Each map set contains 60 instances, resulting in a total of 1800 test instances in this testing benchmark. We assume all agents are homogeneous and share the following parameters: the vehicle’s shape is 3 m × 2 m, with a minimum turning radius of 3 m, and a maximum speed of 1 m/s. All algorithms are assessed using their respective open-source implementations, with most executed on an Intel Xeon Gold 622R CPU at 2.90 GHz in C++, while Fast-ASCO runs on an Intel Core i7-9750H CPU at 2.6 GHz in Matlab.
@@ IV-B Simulation results
TABLE II: Simulation results on random 50 m $\times$ 50 m map.
Method 50m*50m 25 obstacles map with 5,10,15,20,25 agents
Success Rate $\uparrow$ Runtime (s) $\downarrow$ Makespan (s) $\downarrow$
SCC 100.00% 95.00% 81.67% 36.67% 5.00% 0.04 0.42 2.16 7.14 13.39 43.35 48.74 52.21 54.99 54.55
PP 100.00% 95.00% 83.33% 85.00% 51.67% 0.03 0.14 0.36 1.23 2.80 43.35 49.70 53.29 60.03 65.17
FastASCO 90.00% 73.33% 70.00% 65.00% 51.67% 2.46 9.86 24.12 50.64 99.00 56.12 61.58 67.87 73.11 74.73
CSDO 95.00% 95.00% 98.33% 96.67% 90.00% 0.04 0.10 0.25 0.90 3.42 48.42 55.32 59.71 67.06 73.77
50m*50m empty map with 5,10,15,20,25 agents
SCC 100.00% 98.33% 98.33% 70.00% 10.00% 0.03 0.40 1.70 7.19 8.63 41.71 48.22 51.36 52.25 56.34
PP 100.00% 98.33% 96.67% 90.00% 73.33% 0.02 0.09 0.23 0.58 1.28 41.66 48.52 51.75 54.95 59.48
FastASCO 98.33% 96.67% 88.33% 93.33% 85.00% 1.44 7.09 16.52 31.89 53.85 52.25 59.88 63.26 66.26 68.82
CSDO 100.00% 100.00% 98.33% 98.33% 96.67% 0.01 0.03 0.08 0.26 0.83 44.75 52.53 58.31 62.25 66.46
@@
The performance of CSDO is evaluated through a comparative analysis with various MVTP algorithms, focusing on success rate, runtime, and solution quality (makespan). A general time limit of 20 s is applied, except for Fast-ASCO, which is allowed 200 s due to Matlab implementation. Figure 7 presents the results for both map sizes, while Table II displays results for a map size of 50 m $\times$ 50 m.
@@
Seq-CL-CBS (SCC) [2] is a grid search-based method and a prioritized planning version of CL-CBS for large-scale problems. CL-CBS, leveraging the optimal MAPF algorithm Conflict based Search (CBS), forms the basis for SCC, which organizes agents into $k_{b}$ groups for sequential planning with CL-CBS; each subsequent group views the prior as dynamic obstacles. With $k_{b}=2$, SCC performs well with small numbers of agents and generally achieves the best makespan. However, its computation time increases exponentially with large-scale problems due to CBS. In sum, SCC performs worse than CSDO in terms of success rate and runtime.
Prioritized Planning (PP) is a grid search-based prioritized planning method. It randomly assigns each agent an order and plans their trajectories sequentially based on this order. It is implemented by setting $k_{b}=M$ using SCC. In theory, PP is not a complete or optimal method and may perform poorly if an inappropriate order is chosen. In our simulations, PP scales better than SCC, and the solution quality does not deteriorate significantly. However, compared to CSDO, PP exhibits poorer performance in terms of success rate and runtime. CSDO relies on PBS, which searches various partial order sets and has higher completeness. Additionally, CSDO dynamically adjusts its priority during planning to optimize performance. Thus, in crowded scenarios, CSDO may run faster than PP. Regarding solution quality, PP may outperform CSDO, as CSDO prioritizes success rate over makespan optimization.
@@
Fast ASCO [13], an advanced ASCO variant [1], excels in constraint reduction, offering an optimal solution despite slower performance in Matlab. It optimizes for travel time and comfort but has longer makespan due to a detailed kinematics model. Despite this, it achieves superior success rates in large-agent scenarios, outperforming near-optimal methods like SCC for groups of 20 and 25 agents in a 50 m $\times$ 50 m map.
CSDO has the best success rate and computation time, benefiting from PBS’s efficiency and its hierarchical framework. In scenarios with few agents, most methods achieve success rates close to or at 100%. However, as the number of agents increases from 5 to 25 in dense scenarios with 25 obstacles, CSDO maintains a 90% success rate, while existing algorithms begin breaking down, dropping to 51.67%. Moreover, CSDO operates under 1s for 20 agents, significantly outperforming other algorithms.
IV-C CSDO Limitation and Analysis
TABLE III: CSDO analysis on 100m $\times$ 100m map
Number of agents 30 50 70 90
Success rate 96.67% 90.00% 11.67% 1.67%
Search success rate 100% 90.00% 11.67% 1.67%
Centralized Search time(s) 0.87 7.46 9.73 24.72
DQP time(s) 0.06 0.10 0.20 0.28
CSDO is an efficient MVTP algorithm with a high success rate, but it is not an optimal or complete algorithm due to constraints of PBS. As shown in Table III, the success rate is slightly lower than the search rate, indicating that failures are primarily caused by centralized searching. Meanwhile, it’s clear that the search time remains the primary component of the overall runtime. While the simulation demonstrate the efficiency and high success rate of CSDO, MAPF is an NP-hard problem. In simulated crowded scenarios, CSDO may still fail to solve due to timeouts.
@@ V Conclusion
This work introduces CSDO, an efficient algorithm for large-scale multi-vehicle trajectory planning, leveraging a combination of centralized priority-based searching and decentralized optimization. By employing MAPF algorithms, CSDO efficiently identifies feasible solutions to the MVTP optimization challenge. It also accelerates the resolution of MVTP problems efficiently, by decomposing the joint optimization problem into a series of smaller, manageable ones that can be processed in parallel, enabling real-time processing near potential solutions. Through an extensive set of experiments, we demonstrate that CSDO efficiently discovers solutions within a limited time compared to other methods, without significant loss in solution quality, and supports collision-free trajectory planning in large-scale, high-density scenarios.In the future, we will try to extend our algorithm to real-vehicle test environments with dynamic and intensive traffic participants.

### v2

@@ IV-A Simulation Settings
The benchmark consists of various map sets with different sizes (50 m , 100 m), varying numbers of agents (5 to 100), and includes both random and room-like maps. Each map set contains 60 instances, resulting in a total of 2100 test instances in this testing benchmark. We assume all agents are homogeneous and share the following parameters: the vehicle’s shape is 3 m × 2 m, and it has a maximum speed of 1 m/s. All algorithms are assessed using their respective open-source implementations, with most executed on an Intel Xeon Gold 622R CPU at 2.90 GHz in C++, while Fast-ASCO runs on an Intel Core i7-9750H CPU at 2.6 GHz in Matlab.
@@ IV-B Simulation Results
TABLE II: Simulation results on random room-like 100 m $\times$ 100 m map.
Method Success Rate $\uparrow$ Runtime (s) $\downarrow$ Makespan (s) $\downarrow$
# Agents 10 20 30 40 50 10 20 30 40 50 10 20 30 40 50
SCC 93.33% 33.33% 0 0 0 4.43 8.10 - - - 133.02 115.01 - - -
PP 100.00% 91.67% 61.67% 18.33% 6.67% 2.11 7.34 11.02 13.19 13.53 135.41 142.00 131.73 127.35 111.80
FastASCO 30.00% 25% 25% 5% 5% 7.42 35.05 89.28 154.61 378.39 93.65 111.3 110.83 111.64 112.94
CSDO 100.00% 98.33% 80.00% 60.00% 33.33% 1.19 3.64 5.49 7.86 9.19 171.78 187.37 177.71 181.57 178.53
@@
The performance of CSDO is evaluated through a comparative analysis with various MVTP algorithms, focusing on success rate, runtime, and solution quality (makespan). A general time limit of 20 s is applied, except for Fast-ASCO, which is allowed 200 s due to Matlab implementation. Figure 8 presents the results for both map types, while Table II displays results for a map size of 50 m $\times$ 50 m.
@@
Seq-CL-CBS (SCC) [3] is a grid search-based method and a prioritized planning version of the optimal MVTP method, CL-CBS, for large-scale problems. CL-CBS, leveraging the optimal MAPF algorithm Conflict based Search (CBS), forms the basis for SCC, which organizes agents into several groups for sequential planning with CL-CBS; each subsequent group views the prior as dynamic obstacles.
Prioritized Planning (PP) is a grid search-based prioritized planning method. It randomly assigns each agent an order and plans their trajectories sequentially based on this order. In theory, PP is not a complete or optimal method and may perform poorly if an inappropriate order is chosen.
@@
Fast ASCO [15], an advanced ASCO variant [1], excels in constraint reduction, offering an optimal solution despite the high runtime. In addition, it optimizes for both travel time and comfort, and has a detailed kinematics model.
In our simulation, CSDO achieves the best success rate and runtime in general, whether on random maps or room-like maps, benefiting from PBS’s efficiency and its hierarchical framework. Though SCC enhances scalability through sequential planning, its computation time still increases exponentially with large-scale problems due to CBS; PP scales better than SCC, and the solution quality does not deteriorate significantly. However, compared to CSDO, PP exhibits poorer performance in terms of success rate and runtime. CSDO relies on PBS, which searches various partial order sets and has higher completeness. Regarding solution quality, PP may outperform CSDO, as CSDO prioritizes success rate over makespan optimization; As an optimal algorithm, Fast ASCO achieves superior success rates in large scale, outperforming near-optimal methods like SCC for groups of 20 and 25 agents in a 50 m square map. It achieves a longer makespan due to the optimization for comfort and the detailed kinematics model.
IV-C Ablation Study and Limitation Analysis
# Agents 30 50 70 90 SRa: CSDO 96.67% 98.33% 35.00% 5.00% SR: CSDO w/od DOe 0 0 0 0 SR: CSDO w/o warm start 93.33% 95.00% 11.67% 0 FRb: DO failure ratef 1.66% 0 1.66% 0 RTc: CSDO (s) 0.63 3.42 8.99 6.62 RT: CS w/o DO (s) 0.49 3.24 8.74 6.43 RT: CSDO w/o warm start (s) 1.53 9.78 14.72 - a SR: Success Rate; b FR: Failure Rate; c RT: Runtime; d w/o: without; e CSDO w/o DO: Only use centralized search. The success rate means the initial guess has no collision and can be seem as a feasible solution; f DO failure rate: the percentage of cases where decentralized optimization fails to find one feasible solution despite a successful initial guess from centralized searching;
TABLE III: Ablation Study on 100m $\times$ 100m random map
Ablation experiments are conducted to validate the effectiveness of Decentralized Optimization (DO) and a warm start. Without DO, nearly all initial guesses exhibit minor collisions and are deemed infeasible. The runtime of centralized searching dominates within CSDO. Furthermore, the DO failure rate is less than 2%, indicating that the primary completeness loss in CSDO is attributed to PBS. Thus, the necessity, speed, and effectiveness of DO are validated. The warm start technique contributes to varying degrees of improvement in both success rate and runtime metrics.
In summary, regarding limitations, the PBS can be P-complete and can be modified to be P-optimal. For the DO, the completeness drop is less than 2% in simulations.
IV-D Experimental Setup and Results
Refer to caption Refer to caption
(a) (b)
Figure 9: Real world experiments and results on 15 m square map. (a) Experiment platform. (b) The vehicle real trajectories, CSDO planned trajectories and the optimal trajectories planned by CL-CBS on 15 m map.
TABLE IV: Real-world experiment results
Method 15 m $\times$ 15 m map 20 m $\times$ 20 m map
$\tau_{f}$ (s)$\downarrow$ Runtime (s)$\downarrow$ $\tau_{f}$ (s)$\downarrow$ Runtime (s)$\downarrow$
CL-CBS 17.4 10.075 15.5 0.332
CSDO 20.3 0.623 16.8 0.014
Experiments are conducted with 3 forward-only 1.9 m $\times$ 1.3 m vehicles in 15 m and 20 m square areas. The vehicles are positioned using differential GNSS. The map is pre-set and known. The trajectories are calculated on a typical laptop and then transmitted to the agents via WiFi. The trajectory tracking controller operates at 10 Hz, utilizing longitudinal PID control and lateral Pure Pursuit. One scenario and results are shown in Fig. 9. As in Table IV, CSDO achieves similar solution quality while the runtime is greatly reduced.
@@ V Conclusion
This work introduces CSDO, an efficient algorithm for large-scale multi-vehicle trajectory planning, leveraging a combination of centralized priority-based searching and decentralized optimization. Through an extensive set of experiments, we demonstrate that CSDO efficiently discovers solutions within a limited time compared to other methods, without significant loss in solution quality, especially in large-scale, high-density scenarios. In the future, we will try to strike a better balance between solution quality and runtime with SOTA MAPF algorithms, generate robust solutions allowing for tracking errors, and extend our algorithm with dynamic and intensive traffic participants.